prove that cos 2 A+cos 2 B+cos 2 C=1-2cosA.cosB.cocC please solve step by step?

prove that cos2A+cos2B+cos2C=1-2cosA.cosB.cocC please solve step by step?

Grade:11

1 Answers

PASUPULETI GURU MAHESH
8 years ago
cos2A+cos2B+1-sin2ccos2A+cos2B+cos2C=
cos2b+1 =cos(a+c).cos(a-c)+1+cos2b we know a+b+c=1800 cos(a+c)=-cosb =-cosb.cos(a-c)+
                           =cosb(-cos(a-c)-cos(a+c))+1
                          =-cosb(2cosa.cosc)+1
                           =1-2cosa.cosb.cosc
there forecos2A+cos2B+cos2C=1-2cosA.cosB.cocC proved